OpenStudy (anonymous):

factor x^4-x^2+3 = 0 ?

hartnn (hartnn):

let x^2=y so x^4=(x^2)^2=y^2 so now your equation is y^2-y+3=0 can u solve/factor this?

OpenStudy (anonymous):

Dont think it can be factored

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

getting imaginary roots...

OpenStudy (anonymous):

Yep...... we get imaginery roots

hartnn (hartnn):

u have options or something?